# Dynamic Multi-Hop HARQ (DMH-HARQ): Reliable and Open Latency-Constrained Wireless Transport Network

### v3

@@ Abstract
The extreme requirements for high reliability and low latency in the upcoming Sixth Generation (6G) wireless networks are challenging the design of multi-hop wireless transport networks. Inspired by the advent of the virtualization concept in the wireless networks design and openness paradigm as fostered by the O-RAN Alliance, we target a revolutionary resource allocation scheme to improve the overall transmission efficiency.
In this paper, we investigate the problem of multi-hop decode-and-forward (DF) relaying in the finite blocklength (FBL) regime, and propose a DMH-HARQ scheme, which maximizes the end-to-end (E2E) communication reliability in the wireless transport network. We also propose an integer dynamic programming (DP) algorithm to efficiently solve the optimal DMH-HARQ strategy. Constrained within a certain time frame to accomplish E2E transmission, our proposed approach is proven to outperform the conventional listening-based cooperative ARQ, as well as any static HARQ strategy, regarding the E2E reliability. It is applicable without dependence on special delay constraint, and is particularly competitive for long-distance transport network with many hops.
@@ I Introduction
Wireless systems are conceived as a reliable and efficient means to communicate. In particular, with the last generation of wireless networks (5G) unprecedented performance have been achieved while preserving privacy and reliability. In this context, the utilization of relay nodes in wireless communications has proven to be an effective method for combating unexpected network dynamics, such as channel fading, and, in turn, extending coverage [1, 2]. However, with the emergence of novel mobile applications and their increasing demand for exceptional reliability and minimal latency, namely ultra-reliable low-latency communications (URLLC), the feasibility of employing relaying networks in such scenarios has become a challenge [3]. The usage of short codes to meet stringent latency requirements not only significantly reduces the benefits of relaying due to higher error rates [4], but also conventional automatic repeat request (ARQ) solutions struggle to effectively address the diminished link-level reliability as they typically do with longer codes [5].
In parallel, the next generation of wireless networks will power a revolution: network virtualization with a specific focus on radio access networks, a.k.a. virtualized radio access networks (vRAN) [6]. It will hold the promise of substantial cost savings in both capital and operating expenses, while also offering high flexibility and openness to foster innovation and competition. This transformative technology serves as the ultimate milestone as we progress towards the realm of 6G. Nevertheless, extending virtualization to the radio access domain presents a range of challenges that have garnered attention from various initiatives and organizations. Among those, the Open-Radio Access Network (O-RAN) alliance has emerged as a particularly influential force, gaining significant traction and support [7]. O-RAN, a carrier-led initiative, aims to define the next generation of (v)RANs for multi-vendor deployments with a primary objective to disrupt the vRAN ecosystem by dismantling vendor lock-ins and creating an open market that has traditionally been dominated by a small group of players: If O-RAN achieves its goals, it has the potential to unleash an unparalleled level of innovation by reducing the barriers to entry for new competitors [8].
However, the practical deployment of O-RAN in future Sixth Generation (6G) system is facing various emerging challenges, among which there is the essential introduction of wireless transport network to replace the fiber solution. The demand for wireless transport network in 6G is twofold. On the one hand, to support higher throughput and enable accurate positioning, new spectrum at higher frequencies ranging from millimeter-wave (mmWave) to terahertz (THz) will certainly be exploited. The dependence on line-of-sight (LoS) link of these frequencies, however, is implying limited coverage of every individual base station (BS), and inevitably leading to an ultra-dense deployment of small cells [9]. The conventional fiber technology, unfortunately, fails to serve as a reasonably cost-efficient solution for transport networks in this context [9, 10]. On the other hand, aiming at a ubiquitous three-dimensional (3D) coverage, 6G is supposed to deploy non-terrestrial infrastructures such as unmanned aerial vehicles (UAVs), high altitude platforms (HAPs), and satellites, as its serving nodes [11]. Wireless transport network plays here the role of a fundamental enabling technology for the non-terrestrial network (NTN).
In the 6G O-RAN transport network, multi-hop relay mechanisms are integral to the system architecture. User data and control signaling, irrespective of their origin – whether from user equipment (UE) in the uplink (UL) or destined for UE in the downlink (DL)– undergo a sequential relay process through the network’s architecture. The relaying can be executed by a series of network components, including the radio hardware unit (RU), the distributed unit (DU), the central unit (CU), and eventually multiple relay nodes in the case of a long-distance backhaul. Depending on the network functions involved in the communication session and their placement in the network, either part or all of these components can be involved in the relay chain.
Implementing this multi-hop transport network over wireless channels, especially in the context of 6G O-RAN, is challenging conventional radio interface protocols in two aspects. First, compared to classical fiber solutions, wireless technologies are not only suffering from higher propagation loss, exposed to stronger noise and interference, but also more strictly regulated regarding the transmission power. These facts jointly lead to a low signal-to-interference-and-noise ratio (SINR), which challenges the reliability of the radio link, especially when the latency constraint is stringent. Concerning the ambition of 6G for ubiquitous coverage, it will not be a rare case to have long relay chains with many hops in the transport network, which further escalates the conflict between the 6G demands for reliability and latency. To fulfill these requirements, novel methods will be essential to push the transmission performance beyond the state of the art. Second, supporting flexible functional splitting, the O-RAN architecture is also implying a dynamic setup of the relay chain upon the specific choice among different splitting options. Meanwhile, the hybrid 6G architecture consisting of both terrestrial and non-terrestrial infrastructure increases the diversity of single-hop channel features in different deployment scenarios. The high dynamics and diversity of the multi-hop transport network are challenging conventional radio protocols based on fixed radio frame design and static resource scheduling.
Motivated by the need to enhance openness and reliability to end-to-end (E2E) links while fulfilling strict time constraints imposed by URLLC applications in a multi-hop transport netork, we draw inspiration from the concept of dynamically reallocating the remaining radio resources between repeats and successor hops based on instantaneous transmission results. Therefore, in this paper we pioneer the design of a novel dynamic hybrid ARQ (HARQ) scheme for multi-hop relaying in the finite blocklength (FBL) regime. We conduct a comprehensive analysis of its performance and propose a dynamic programming (DP) algorithm for optimizing its schedule. By exploring this approach, we aim to address the challenges faced by traditional relaying networks and improve the overall reliability of the O-RAN-inspired wireless transport network. Our research focuses on achieving enhanced E2E reliability through dynamic resource allocation, taking into account the specific requirements and constraints of modern mobile systems operating in high-demand scenarios.
@@ II Related Work
Over the past few years, the emerging wireless use scenario of URLLCin O-RAN is challenging off-the-shelf multi-hop relay. The extreme requirement of low E2E latency in O-RAN URLLC is leading to even more stringent time constraint over each individual hop [19, 20], which is barely achievable with conventional technologies. To meet such a stringent latency requirement, URLLC transmissions are likely carried out with short packet, i.e., the codewords cannot considered as infinite long. With such so-called FBL codes, many theories and propositions in classical information theory fail to apply since the Shannon limit is no more asymptotically achievable. To accurately characterize the relationship between achievable data rate and reliability in the FBL regime, Polyanskiy et al. derived in [21] a closed-form expression for the single-hop transmission in additive white Gaussian noise (AWGN) channels. Later on, this FBL characterization was generalized, extended into Gilbert-Elliot channels [22] and quasi-static flat-fading channels [23]. Recently, the FBL model in random access channel has also been analyzed [24]. On the basis of those results, the performance FBL communication systems has been investigated in context of various wireless networking technologies, such as multi-input multi-output (MIMO) [25], orthogonal frequency-division multiple access (OFDMA) [26], non-orthogonal multiple access (NOMA) [27] and multi-access edge computing (MEC) [28].
@@ III-A Multi-Hop DF and DMH-HARQ
Now consider a multi-hop relaying chain with $I+1$ nodes over $I$ hops, each applying the DF strategy based on simple HARQ without combining. We assume that all channels of the $I$ hops are i.i.d., undergoing Gaussian noise and Rician fading. We assume that every decoding process, together the associated feedback of acknowledgement (ACK) or negative acknowledgement (NACK) to the predecessor node for the HARQ procedure, makes the same delay of $T_{\mathrm{l}}$. The overall transmission to a maximal E2E delay is limited to $T_{\mathrm{max}}$. For simplification of the analysis, we consider for every node the same symbol rate $f_{\mathrm{s}}$, modulation order $M$ (and therefore the bit-time length $T_{\mathrm{b}}$=$\frac{1}{Mf_{\mathrm{s}}}$), transmission power $P$, and noise power $N$. The transmitting node of each hop is capable of flexibly adjust its channel coding rate in every (re-)transmission attempt independently from others. We consider in this study a block fading scenario, where the channel $h_{i}$ of each hop $i\in\mathcal{I}=\{1,2,\dots I\}$ remains consistent over the entire time frame of $T_{\mathrm{max}}$. Each transmitting node of hop $i$ possesses the perfect CSI of $h_{i}$, while knowing only its position in the relaying chain and the statistical CSI of all other hops.
Here we consider a problem of dynamic radio resource allocation among different HARQ slots over the $I$-hop DF chain. On every hop, the transmitting node can individually adjust its coding rate for the next (re)transmission. Since the modulation order and symbol rate are fixed, this will determine the time span of the transmission slot, which cannot exceed the remaining time of the entire transmission frame. If the message is successfully decoded, the successor node will take the turn and further forward the message, unless it is the sink node, i.e. the $(I+1)^{\mathrm{th}}$. If the decoding fails, the receiving node will trigger the HARQ procedure by sending back a NACK to the transmitting node, which then keeps re-transmitting the message until either a successful reception or a timeout, i.e. when the remaining time falls insufficient to support any further effective transmission. In every individual repeat, the node is able to readjust its coding rate. We consider a timer embedded in the header of every message, so that every node is aware of the time remaining from the $T_{\mathrm{max}}$ frame. We name this proposed approach as Dynamic Multi-Hop HARQ (DMH-HARQ), which is briefly illustrated in Fig. 1.
Figure 1: Illustration of a DMH-HARQ schedule for 4-hop DF chain, where the remaining (reserved) blocklength of a frame is dynamically assigned regarding the decoding result: either for next hops upon success, or for retransmission on the same hop upon failure.
@@ III-B Binary Schedule Tree
Each node $s$ in the tree representing a certain system state, which can be defined as a three tuple $s=(t_{s},i_{s},k_{s},\tau_{s})$ where $t_{s}\in\mathbb{R}^{+}$ is the remaining time, $i_{s}\in\mathbb{N}$ is the remaining hops to transmit the message, $k_{s}\in\mathbb{N}$ is the number of failed HARQ attempts on the current hop, and $\tau_{s}$ the accumulated time spent on these $k_{s}$ unsuccessful transmissions (excluding the feedback losses). Now let $n_{s}\triangleq n(t_{s},i_{s},k_{s},\tau_{s})\in\mathbb{N}$ denote the the blocklength to be allocated, according to the policy $n(\cdot)$, to the current HARQ attempt on the current hop, i.e. the $(k_{s}+1)$th HARQ attempt on the $i_{s}$th last hop, when there is $t_{s}$ time remaining. Thus, let $\mathfrak{T}_{s}$ denote a schedule tree with $S(\mathfrak{T}_{s})=\{s\}$ and $\mathcal{S}$ the space of all feasible states, for all $s\in\mathcal{S}$, on the one hand we have for the left branch
@@ III-C Optimization Problem
To measure the effectiveness of a DMH-HARQ schedule, we define a utility $\xi(\mathfrak{T}_{s})$ as the chance of successfully forwarding the message over the next $i_{s}$ hops within a time limit of $t_{s}$, then from the recursive structure (12) of $\mathfrak{T}_{s}$, we have
@@
$n_{\mathrm{min},i_{s}}=\min\{n\in\mathbb{N}^{+}~{}|~{}\varepsilon_{s}\leqslant\varepsilon_{\mathrm{max}}\}$ (6)
@@ IV-A Error Model, Constraints, and Approximations
Regarding Type I HARQ without combining, according to [21], the PER of an arbitrary HARQ attempt with blocklength $m$ on the $i^{\mathrm{th}}$ hop is
$\epsilon_{i}(m)=Q\left(\sqrt{\frac{m}{V_{i}}}\left(\mathcal{C}_{i}-\frac{d}{m}\right)\ln 2\right),$ (7)
where for AWGN channels $V_{i}=1-\frac{1}{(1+\gamma_{i})^{2}}$ with $\gamma_{i}=\frac{Pg_{i}}{N}=\frac{P\overline{g}}{NL_{\text{f},i}}.$ denoting the SNR on the $i^{\mathrm{th}}$ hop. Here, $g_{i}$ is the channel gain of the $i^{\mathrm{th}}$ hop, which is determined by the mean channel gain $\overline{g}$ and the random fading loss $L_{\text{f},i}$. Thus, for all $s=(t_{s},i_{s},k_{s},\tau_{s})\in\mathcal{S}$, with the blocklength allocation policy $n_{s}$, we have
$\varepsilon_{s}=\epsilon_{I+1-i_{s}}(n_{s})=Q\left(\sqrt{\frac{n_{s}}{V_{I+1-i_{s}}}}\left(\mathcal{C}_{I+1-i_{s}}-\frac{d}{n_{s}}\right)\ln 2\right),$ (8)
@@
Following the common routine in FBL study, we can extend the space of blocklength from integer to real numbers, so as to allow non-integer values of $n_{s}$ for the convenience of analysis (approximate integer solutions can be then obtained by rounding the real-valued optima). Thus, the original problem (III-C) can be replaced by {maxi} n:SR^+ξ(T_s_0)
Recalling (6), the lower-bound $n_{\mathrm{min},i_{s}}$ in constraint (III-C) can be obtained from:
$Q\left(\sqrt{\frac{n_{\mathrm{min},i_{s}}}{V_{I+1-i_{s}}}}\left(\mathcal{C}_{I+1-i_{s}}-\frac{d}{n_{\mathrm{min},i_{s}}}\right)\ln 2\right)=\varepsilon_{\mathrm{max}},$ (9)
which depends on $s$ only regarding $i_{s}$. Meanwhile, the upper-bound $n_{\mathrm{max},s}\triangleq\frac{t_{s}-T_{\mathrm{min}}^{(i_{s}-1)}}{Mf_{\mathrm{s}}}$ is determined by both $t_{s}$ and all $n_{\mathrm{min},j}$ that $11\leqslant j\leqslant i_{s}-$ 1, i.e., by $t_{s}$ and $i_{s}$.

### v4

# DMH-HARQ: Reliable and Open Latency-Constrained Wireless Transport Network

@@ Abstract
The extreme requirements for high reliability and low latency in the upcoming Sixth Generation (6G) wireless networks are challenging the design of multi-hop wireless transport networks. Inspired by the advent of the virtualization concept in the wireless networks design and openness paradigm as fostered by the Open-Radio Access Network (O-RAN) Alliance, we target a revolutionary resource allocation scheme to improve the overall transmission efficiency.
In this paper, we investigate the problem of automatic repeat request (ARQ) in multi-hop decode-and-forward (DF) relaying in the finite blocklength (FBL) regime, and propose a dynamic scheme of multi-hop hybrid ARQ (HARQ), which maximizes the end-to-end (E2E) communication reliability in the wireless transport network. We also propose an integer dynamic programming (DP) algorithm to efficiently solve the optimal Dynamic Multi-Hop HARQ (DMH-HARQ) strategy. Constrained within a certain time frame to accomplish E2E transmission, our proposed approach is proven to outperform the conventional listening-based cooperative ARQ, as well as any static HARQ strategy, regarding the E2E reliability. It is applicable without dependence on special delay constraint, and is particularly competitive for long-distance transport network with many hops.
@@ I Introduction
Ultra-reliable low-latency communications (URLLC) is pushing wireless transport beyond fiber-based backhaul toward multi-hop wireless chains that stitch together terrestrial small cells and non-terrestrial segments within virtualized, Open-Radio Access Network (O-RAN) [6, 7, 8]. In such deployments, fronthaul, midhaul, and backhaul flows may traverse sequences of open radio hardware unit (O-RU), open distributed unit (O-DU), and open centralized unit (O-CU) nodes hosted on O-Cloud and orchestrated by the SMO, with the specific chain length and composition dictated by functional splits and dynamic placement.
The shift to millimeter-wave (mmWave) and (sub-)THz for capacity and positioning tightens line-of-sight constraints and shrinks cell footprints, making dense, relay-rich topologies economically preferable to ubiquitous fiber [9, 10]. At the same time, 3D coverage objectives fold in non-terrestrial network (NTN) elements—unmanned aerial vehicles (UAVs), high altitude platforms (HAPs), and satellites—whose heterogeneous channels and feedback latencies further complicate transport [11]. While relays have long been used to extend coverage and combat fading [1, 2], URLLC alters the operating point: packets are short (finite blocklength), feedback and control overheads are non-negligible, and strict end-to-end (E2E) deadlines mean that each retransmission on one hop directly consumes the time/frequency budget of successor hops. Under these conditions, classical relaying and ARQ/hybrid ARQ (HARQ) designs (tuned for long codes and per-hop optimization) systematically misestimate reliability and cannot coordinate retransmissions and forwarding at the granularity and speed demanded by URLLC [4, 5].
Challenges. The first challenge is reliability composition under finite blocklength (FBL): with short codes, per-hop decoding errors remain non-negligible and compound across hops, making E2E reliability exquisitely sensitive to how retransmissions and forwarding are budgeted within a tight delay. Second, resources are coupled across hops; allocating more symbols to a retransmission on one hop reduces the symbols—and thus reliability—available to successor hops. Third, links at mmWave and (sub-)THz under strict power limits and interference are brittle, while NTN segments (UAVs, HAPs, satellites) introduce heterogeneous channels and feedback delays. Fourth, open and disaggregated O-RAN deployments with flexible functional splits (O-RU/O-DU/O-CU, O-Cloud, SMO) yield relay chains whose length and per-hop characteristics vary with placement and load, challenging fixed frame designs and static schedulers.
Figure 1: A conceptual depiction of the O-RAN architecture is presented. This article concentrates on the areas outlined by the dashed blue box, where the proposed transmission scheme enhances E2E communication reliability.
Limitations of existing methods. Classical relaying assumes long codes and asymptotic coding gains, which do not hold in the FBL regime of short packets [4, 5]. Standard ARQ/HARQ mechanisms are tuned to per-hop reliability or subframe-level allocation and do not perform outcome-conditioned reallocation under a global E2E deadline, thereby neglecting the inter-hop coupling central to URLLC. Fixed retransmission counts, timers, or frame partitions cannot adapt to instantaneous decode outcomes, hop heterogeneity, or O-RAN-driven topology changes. Even listening-based cooperative automatic repeat request (ARQ) [12] typically moves resources at the frame/subframe timescale: when packets are short, symbol-level decisions become pivotal yet are unsupported.
Contributions. We propose a dynamic HARQ scheme for multi-hop relaying in the FBL regime that reallocates—at channel-use (symbol) granularity—the remaining radio budget between the current hop’s repeats and successor hops, conditioned on instantaneous decode outcomes. The design is O-RAN-compatible (Fig. 1) and targets open, disaggregated transport. Our key insight is to treat retransmissions and forwarding as a single coupled E2E optimization rather than per-hop tuning, updating the residual budget after each (re)transmission using posterior success probabilities under FBL. We formalize the scheduler and provide a dynamic programming (DP) algorithm that optimizes the sequence of (re)transmissions across hops under latency and reliability targets.
In summary, our contributions are as follows: $i$) a dynamic, symbol-level HARQ framework that explicitly couples retransmission and forwarding under a strict E2E deadline, $ii$) an FBL-based analysis that quantifies reliability composition across hops, and iiiiii) a DP-based optimal scheduler with implementable policies aligned with O-RAN disaggregation. Compared to static HARQ, subframe-level listening-based cooperative ARQ [12], and long-code-inspired relaying [1, 2], our approach delivers improved reliability–latency trade-offs for URLLC over O-RAN-inspired wireless transport.
@@ II Related Work
Over the past few years, the emerging wireless use scenario of URLLCin O-RAN is challenging off-the-shelf multi-hop relay. The extreme requirement of low end-to-end (E2E) latency in O-RAN URLLC is leading to even more stringent time constraint over each individual hop [22, 23], which is barely achievable with conventional technologies. To meet such a stringent latency requirement, URLLC transmissions are likely carried out with short packet, i.e., the codewords cannot considered as infinite long. With such so-called FBL codes, many theories and propositions in classical information theory fail to apply since the Shannon limit is no more asymptotically achievable. To accurately characterize the relationship between achievable data rate and reliability in the FBL regime, Polyanskiy et al. derived in [24] a closed-form expression for the single-hop transmission in additive white Gaussian noise (AWGN) channels. Later on, this FBL characterization was generalized, extended into Gilbert-Elliot channels [25] and quasi-static flat-fading channels [26]. Recently, the FBL model in random access channel has also been analyzed [27]. On the basis of those results, the performance FBL communication systems has been investigated in context of various wireless networking technologies, such as multi-input multi-output (MIMO) [28], orthogonal frequency-division multiple access (OFDMA) [29], non-orthogonal multiple access (NOMA) [30] and multi-access edge computing (MEC) [31].
@@ III-A Multi-Hop DF and DMH-HARQ
Now consider a multi-hop relaying chain with $I+1$ nodes over $I$ hops, each applying the DF strategy based on simple HARQ without combining. We assume that all channels of the $I$ hops are i.i.d., undergoing Gaussian noise and Rician fading. We assume that every decoding process, together the associated feedback of acknowledgement (ACK) or negative acknowledgement (NACK) to the predecessor node for the HARQ procedure, makes the same delay of $T_{\mathrm{l}}$. The overall transmission to a maximal E2E delay is limited to $T_{\mathrm{max}}$. For simplification of the analysis, we consider for every node the same symbol rate $f_{\mathrm{s}}$, modulation order $Q_{\mathrm{m}}$ (and therefore the bit-time length $T_{\mathrm{b}}$$\frac{1}{{Q_{\mathrm{m}}}f_{\mathrm{s}}}$), transmission power $P$, and noise power $N$. The transmitting node of each hop is capable of flexibly adjust its channel coding rate in every (re-)transmission attempt independently from others. We consider in this study a block fading scenario, where the channel $h_{i}$ of each hop $i\in\mathcal{I}=\{1,2,\dots I\}$ remains consistent over the entire time frame of $T_{\mathrm{max}}$. Each transmitting node of hop $i$ possesses the perfect CSI of $h_{i}$, while knowing only its position in the relaying chain and the statistical CSI of all other hops.
Here we consider a problem of dynamic radio resource allocation among different HARQ slots over the $I$-hop DF chain. On every hop, the transmitting node can individually adjust its coding rate for the next (re)transmission. Since the modulation order and symbol rate are fixed, this will determine the time span of the transmission slot, which cannot exceed the remaining time of the entire transmission frame. If the message is successfully decoded, the successor node will take the turn and further forward the message, unless it is the sink node, i.e. the $(I+1)^{\mathrm{th}}$. If the decoding fails, the receiving node will trigger the HARQ procedure by sending back a NACK to the transmitting node, which then keeps re-transmitting the message until either a successful reception or a timeout, i.e. when the remaining time falls insufficient to support any further effective transmission. In every individual repeat, the node is able to readjust its coding rate. We consider a timer embedded in the header of every message, so that every node is aware of the time remaining from the $T_{\mathrm{max}}$ frame. We name this proposed approach as Dynamic Multi-Hop HARQ (DMH-HARQ), which is briefly illustrated in Fig. 2 in context of an example 4-hop DF chain.
Figure 2: Part of a DMH-HARQ schedule for 4-hop DF chain, beginning with the initial state $s_{0}=(T_{\mathrm{max}},4,0,0)$. At each (re)transmission attempt, an optimal blocklength is calculated for consumption, and a fixed time for decoding and feedback is required. Upon the decoding result, the remaining (reserved) blocklength is dynamically assigned regarding the decoding result: either for next hops upon success (ACK), or for retransmission on the same hop upon failure (NACK). The states $s_{0}$, $s_{1}$, and $s_{2}$ are illustrated with time frame details.
@@ III-B Binary Schedule Tree
Each node $s$ in the tree representing a certain system state, which can be defined as a 4-tuple $s=(t_{s},i_{s},k_{s},\tau_{s})$ where $t_{s}\in\mathbb{R}^{+}$ is the remaining time, $i_{s}\in\mathbb{N}$ is the remaining hops to transmit the message, $k_{s}\in\mathbb{N}$ is the number of failed HARQ attempts on the current hop, and $\tau_{s}$ the accumulated time spent on these $k_{s}$ unsuccessful transmissions (excluding the feedback losses). This system state concerns various aspects of the relaying session, involving the remaining resource, the transmissions to accomplish, and partial history of previous transmissions. Take the example in Fig. 2, starting with the initial state $s_{0}=(T_{\mathrm{max}},4,0,0)$, where the schedule assigns $n_{s_{0}}$ bits for the first transmission, reserving therewith $r_{s_{0}}=T_{\mathrm{max}}-n_{s_{0}}T_{\mathrm{b}}-T_{\mathrm{l}}$ for future (re)transmissions. Upon the first transmission result, two possible succeding states can be $s_{1}=(r_{s_{0}},4,1,n_{s_{0}}T_{\mathrm{b}})$ and $s_{2}=(r_{s_{0}},3,0,0)$, respectively. Now let $n_{s}\triangleq n(t_{s},i_{s},k_{s},\tau_{s})\in\mathbb{N}$ denote the the blocklength to be allocated at state $s$, according to the policy $n(\cdot)$, to the current HARQ attempt on the current hop, i.e. the $(k_{s}+1)$th HARQ attempt on the $(I-i_{s}+1)^{\mathrm{th}}$ hop, when there is $t_{s}$ time remaining. Thus, let $\mathfrak{T}_{s}$ denote a schedule tree with $S(\mathfrak{T}_{s})=\{s\}$ and $\mathcal{S}$ the space of all feasible states, for all $s\in\mathcal{S}$, on the one hand we have for the left branch
@@ III-C Optimization Problem
To measure the effectiveness of a DMH-HARQ schedule, we define a utility $\xi(\mathfrak{T}_{s})$ as the chance of successfully forwarding the message over the next $i_{s}$ hops within a time limit of $t_{s}$, then from the recursive structure (1)–(2) of $\mathfrak{T}_{s}$, we have
@@
Thus, given the time frame length $T_{\mathrm{max}}$ and total number of hops $I$ specified, the reliability of DMH-HARQ can be optimized regarding the blocklength allocation policy $n$:
where
@@ IV-A Error Model, Constraints, and Approximations
Regarding Type I HARQ without combining, according to [24], the PER of an arbitrary HARQ attempt with payload length $d$ and blocklength $m$ on the $i^{\mathrm{th}}$ hop is
$\epsilon_{i}(m)=Q\left(\sqrt{\frac{m}{V_{i}}}\left(\mathcal{C}_{i}-\frac{d}{m}\right)\ln 2\right),$ (6h)
where for AWGN channels $V_{i}=1-\frac{1}{(1+\gamma_{i})^{2}}$, with $\gamma_{i}=\frac{Pg_{i}}{N}=\frac{P\overline{g}}{NL_{\text{f},i}}$ denoting the SNR on the $i^{\mathrm{th}}$ hop. Here, $g_{i}$ is the channel gain of the $i^{\mathrm{th}}$ hop, which is determined by the mean channel gain $\overline{g}$ and the random fading loss $L_{\text{f},i}$. Thus, for all $s=(t_{s},i_{s},k_{s},\tau_{s})\in\mathcal{S}$, with the blocklength allocation policy $n_{s}$, we have
$\varepsilon_{s}=\epsilon_{I+1-i_{s}}(n_{s})=Q\left(\sqrt{\frac{n_{s}}{V_{I+1-i_{s}}}}\left(\mathcal{C}_{I+1-i_{s}}-\frac{d}{n_{s}}\right)\ln 2\right),$ (6i)
@@
Following the common routine in FBL study, we can extend the space of blocklength from integer to real numbers, so as to allow non-integer values of $n_{s}$ for the convenience of analysis (approximate integer solutions can be then obtained by rounding the real-valued optima). Thus, the original problem (6) can be replaced by
[…]
Recalling (6g), the lower-bound $n_{\mathrm{min},i_{s}}$ in constraint (6a) can be obtained from:
$Q\left(\sqrt{\frac{n_{\mathrm{min},i_{s}}}{V_{I+1-i_{s}}}}\left(\mathcal{C}_{I+1-i_{s}}-\frac{d}{n_{\mathrm{min},i_{s}}}\right)\ln 2\right)=\varepsilon_{\mathrm{max}},$ (6k)
which depends on $s$ only regarding $i_{s}$. Meanwhile, the upper-bound $n_{\mathrm{max},s}\triangleq\frac{t_{s}-T_{\mathrm{min}}^{(i_{s}-1)}}{Mf_{\mathrm{s}}}$ is determined by both $t_{s}$ and all $n_{\mathrm{min},j}$ that $1\leqslant j\leqslant i_{s}-1$, i.e., by $t_{s}$ and $i_{s}$.